34 49 34 36 44 54 44 50 46 43 34 43 31

45 39 34 47 44 25 38 54 53 47 37 40

  1. How many classes would you recommend?
  2. What class interval would you suggest? (Round up your answer to the next whole number.)

Two thousand frequent business travelers are asked which midwestern city they prefer: Indianapolis, Saint Louis, Chicago, or Milwaukee. One hundred and eight liked Indianapolis best, 432 liked Saint Louis, 1258 liked Chicago, and the remainder preferred Milwaukee. Develop a frequency table and a relative frequency table to summarize this information. (Round the Relative Frequency to 3 decimal places.)

   City Frequency Relative frequency

Indianapolis

St. Louis

Chicago

Milwaukee
Answer #1

Problem 1:

Maximum value=54, minimum value=25, range=54-24=29

Take no. of classes=5

Width of class interval=29/5=5.8 approximately 6.

Problem 2

Frequency (f) Relative frequency ((f/2000) 100) City Indianapolis St. Louis Chicago Milwaukee Total 108 432 1258 202 2000(EN) 5.4 % 21.6 % 62.9 % 10.1% 100 %